Use the definition of limit to prove that the sequence {n-1/n} n=1 to infinity is divergent


1
Expert's answer
2021-01-28T05:34:12-0500

Solution


The number "L" is the limit of the sequence "\{a_n\}" if

"(1)\ \epsilon > 0, \ a_n \underset{e}{\simeq } L, for\ n \geq 1."

If such an L exists, we say "\{a_n\}"converges, or is convergent; if not, "\{a_n\}" diverges,or is divergent.


Let "L\u2208R". We claim that the sequence "s_n=\begin{Bmatrix}
n-\frac{1}{n}
\end{Bmatrix}n" does not converge to L. To show this, take "\u03b5= 1". Since "s_n=\begin{Bmatrix}
n-\frac{1}{n}
\end{Bmatrix}n \geq n" , if "n >|L|+ 1" we have that



"|s_n\u2212L|\u2265|s_n|\u2212|L|\u2265n\u2212|L|>1 =\u03b5."


It follows that for any "N\u2208N", if we take "n >max\{|L|+ 1,N\}" then "|sn\u2212L|> \u03b5" . Thus by the definition of the limit "(s_n)" diverges.

The notations for the limit of a sequence in this case are:


"\overset{lim}{n \to \infty} \begin{Bmatrix}
n-\frac{1}{n}
\end{Bmatrix}= \infty"

This sequence "\begin{Bmatrix}
n-\frac{1}{n}
\end{Bmatrix}_{n = 1}" is divergent